Find the x intercept and y intercept of each equation:

y= 3x-6

y= -2x

x+y= 5



1
Expert's answer
2021-01-14T19:59:14-0500

For intersection with OX the y=0, for intersection with OY the x=0. So in the first equation we have

OX

3x-6=0

3x=6

x=2, so the point is (2; 0)


OY

y=3*0-6

y=0-6

y=(-6), so the point is (0; -6)


The same method for two other equations.

In total answer is:

 X    Y

1) (2;0) (0;-6)

2) (0;0)  (0;0)

3) (5;0) (0;5)